# Complementary Advantages: Exploiting Cross-Field Frequency Correlation for NIR-Assisted Image Denoising

### v1

@@ 3.5 Loss Function
The setting of the loss function is similar to DVN, adopting the Charbonnier loss [4], and adding a frequency domain loss [8]. The overall loss function is represented as
$\begin{split}\mathcal{L}=\mathcal{L}_{charbonnier}(\hat{T}_{1},G)+\mathcal{L}_{charbonnier}(\hat{T}_{2},G)\\ +\alpha\cdot\mathcal{L}_{charbonnier}(\mathcal{F}(\hat{T}_{1}),\mathcal{F}(G)),\end{split}$ (6)
where $\hat{T}_{1}$ and $\hat{T}_{2}$ are the outputs of pre-denoising sub-network and whole network, respectively. $G$ is ground truth and $\alpha$ is set as 0.1.
@@
4 Experiments
In this section, we evaluate our method on public datasets and compare it with other NIR-assisted RGB image denoising methods.
@@ 4.1 Experimental Settings
Datasets. We evaluate the NIR-guided RGB image denoising method on the DVD [16] and IVRG [3] datasets. The DVD dataset consists of 307 high-resolution NIR-RGB image pairs, which are converted into 5k image pairs of size $3\times 256\times 256$ for training and 1k image pairs of size $3\times 256\times 256$ for testing. The IVRG dataset contains NIR and RGB image pairs from 9 different scenes, including countryside, fields, forests, indoors, mountains, old buildings, streets, cities, and water. For each scene, as in [39], experiments randomly select 30/8/15 pairs for training/validation/testing.
Parameter settings. The proposed network is trained using the Adam optimizer [11] ($\beta_{1}=0.9$, $\beta_{2}=0.999$) on patches of size $128\times 128$ with a batch size of 16. For the DVD dataset, we simulate low-light conditions by randomly reducing the mean of the original images as in [16] and add Gaussian-Poisson mixed noise with a range of 1 to 16. The model is trained for 80 epochs. The initial learning rate is $2\times 10^{-4}$, and a cosine annealing strategy is used to gradually reduce the learning rate to $1\times 10^{-6}$. For the IVRG dataset, following the settings in [39], it is cropped into 16,374 training patches, and then additive Gaussian noise with random noise levels is used to simulate the target images, considering three noise levels in testing: $\sigma=25$, 50, and 75. The model is trained for 100 epochs with an initial learning rate of $5\times 10^{-4}$. All models are implemented using PyTorch on an NVIDIA RTX 3090 GPU.
@@
Restormer 31.72 0.876 27.67 0.788 24.50 0.722
Uformer 31.97 0.878 27.83 0.793 24.58 0.726
MPRNet 31.89 0.878 27.80 0.794 24.57 0.729
Xformer 31.98 0.879 27.86 0.794 24.61 0.731
CUNet 31.37 0.876 27.78 0.812 24.64 0.764
SANet 31.54 0.885 28.13 0.825 24.81 0.785
MNNet 32.59 0.904 28.66 0.851 25.25 0.808
DVN 32.76 0.906 28.73 0.852 25.28 0.809
NIR-Restormer 32.86 0.906 28.77 0.853 25.30 0.811
@@ 4.2 Experimental Results
Evaluations on the DVN dataset. We first evaluate our method on the DVD dataset, with quantitative results shown in Table 1. It can be observed that our FCENet, and even FCENet-L, outperform all other single-image denoising methods and NIR-assisted denoising methods in terms of PSNR and SSIM with moderate computational and memory costs. The qualitative comparison in Fig. 6 clearly illustrates that our method has a significant advantage over other state-of-the-art methods in noise removal, color preservation, and detail recovery. For single-image denoising methods like Restormer, although they have achieved advanced denoising effects in low-noise scenarios, they still lack effective means to address detail texture loss in high-noise situations. For guided denoising methods, MNNet is unable to effectively handle the significant inconsistencies between NIR and RGB. DVN strengthens consistent high-frequency textures but loses critical differential high-frequency information in NIR and sacrifices the authenticity of low-frequency color; NIR-restormer
Figure 8: Frequency domain dynamic filters on NIR and RGB side at different depths. The first row represents the frequency-domain filters of RGB side, and the second row represents the filters of NIR side.
has good denoising effects but lacks effective utilization of inconsistent information, leading to over-smoothing and color distortion.Our method recovers detailed texture information while more faithfully preserving low-frequency color information in RGB. This proves that our proposed method, through a series of frequency domain learning, can fully exploit useful information from two fields and better solve the fusion problem under the inconsistency of color and structure between NIR and RGB images.
Evaluations on the IVRG dataset. We evaluate our method on the IVRG dataset. Following the previous method, we consider three noise levels in the test, including $\sigma=25,50$, and $75$. Quantitative results are shown in Table 2. It can be observed that compared to the state-of-the-art single-image denoising algorithms, our method significantly improves the performance of image denoising, demonstrating the effectiveness of near-infrared images. Compared to existing NIR-assisted denoising methods, our method also has significant advantages. More visualization results are placed in the supplementary material.
Evaluations on real-world experiment. To further evaluate the performance of our model in real noisy environments, we conduct qualitative experiments on NIR-RGB data pairs in real low-light conditions. As shown in Fig. 7, the introduction of NIR images brings more detailed textures but also a large amount of inconsistency.
@@
DRM 30.15 0.948
DRM + CRM 30.37 0.950
SANet fails to handle inconsistencies, resulting in a significant amount of artifacts, while MNNet, DVN and NIR-Restormer are unable to thoroughly denoise. In contrast, our method produces a smoother and more natural effect, eliminating inconsistencies while introducing more detailed information.
@@ 4.3 Ablation Study
We conduct ablation studies on the DVD dataset using our model. The ablation experiments are evaluated under the condition of $\sigma=4$. We first compare the two mechanisms proposed in this paper with a two-stage U-Net baseline, with the results shown in Table 3. In the U-Net baseline, cross-field images are fused through addition. By deploying FDSM and FEFM, the model achieve gains of 0.51dB and 0.83dB, respectively. Finally, combining all contributions significantly improve the denoising level compared to the baseline.
Effect of FDSM. Using FDSM allows for better extraction of complementary information from NIR and RGB while handling inconsistent regions. In Fig. 8, we show the frequency domain dynamic filter kernels in FDSM.
@@
It can be observed that the filter kernels processing NIR feature maps regularly enhance the learning of mid-to-high frequency information, while those processing RGB feature maps mainly learn low frequency information from the input images. This is consistent with the conclusions drawn from our frequency domain correlation priors. Additionally, we conducted experiments using spatial dynamic convolution in this module. The results in Table 4 ultimately prove the effectiveness of extracting relevant features from the frequency domain in NIR and RGB images and the stability of dynamic convolution in extracting inconsistent features.
Effect of FEFM. Traditional cross-attention fusion modules only strengthen common features but neglect differential features. FEFM module strengthens the learning of common features and enhances the learning of some high-frequency differential information in NIR feature maps. Fig. 9 shows the input NIR and RGB feature maps after frequency domain selection, the fused feature maps after cross-attention fusion, and the fused feature maps after FEFM. Compared to the NIR feature maps, the input RGB feature maps still lack significant high-frequency features. It can be seen that our FEFM strengthens common features while supplementing some differential texture information from NIR. We compare two mechanisms in FEFM, the Common feature Reinforcement Mechanism (CRM) and the Differential feature Reinforcement Mechanism (DRM), with traditional cross-attention fusion modules and direct feature summation. The results in Table 5 prove effectiveness of our method in strengthening common features in frequency domain and modeling differential features.
@@ 5 Conclusion
In this paper, we propose a cross-field frequency correlation exploiting network for NIR-assisted image denoising, termed FCENet. Specifically, we first conduct a series of analyses on the characteristics of NIR and RGB images in frequency domain and obtain frequency correlation prior. Based on the prior, our frequency dynamic selection mechanism dynamically extracts complementary information from cross-field images in the frequency domain, and we develop a frequency exhaustive fusion mechanism to enhance common features and differential high-frequency features of two fields during fusion process. Extensive experiments on public NIR-RGB datasets demonstrate effectiveness of frequency domain learning, and our method achieves better results than state-of-the-art methods.

### v2

@@ 3.5 Loss Function
We adopt the Charbonnier loss [5], and adding the frequency domain loss [9]. The overall loss function is represented as
$\begin{split}\mathcal{L}=\mathcal{L}_{charbonnier}(\mathbf{X}_{1},\mathbf{T})+\mathcal{L}_{charbonnier}(\mathbf{X}_{2},\mathbf{T})\\ +\alpha\cdot\mathcal{L}_{charbonnier}(\mathcal{F}(\mathbf{X}_{2}),\mathcal{F}(\mathbf{T}))\end{split}$ (6)
where $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ are the outputs of pre-denoising sub-network and whole network, respectively. $\mathbf{T}$ is ground truth and $\alpha$ is set as 0.1 [9].
4 Experiments
In this section, we evaluate the proposed method on public datasets and compare it with other NIR-assisted image denoising methods.
@@ 4.1 Experimental Settings
Datasets. We evaluate the NIR-assisted image denoising method on the DVD [18] and IVRG [4] datasets. The DVD dataset consists of 307 high-resolution NIR-RGB image pairs, which are converted into 5000 image pairs of size $3\times 256\times 256$ for training and 1000 image pairs of size $3\times 256\times 256$ for testing. The IVRG dataset contains NIR and RGB image pairs from 9 different scenes. For each scene, as in [44], experiments randomly select 30/8/15 pairs for training/validation/testing.
Parameter settings. The proposed network is implemented with Pytorch and trained using the Adam optimizer [20] ($\beta_{1}=0.9$, $\beta_{2}=0.999$) on patches of size $128\times 128$ with a batch size of 16. For the DVD dataset, we simulate low-light conditions by randomly reducing the mean of the original images as in [18] and add Gaussian-Poisson mixed noise with a range of 1 to 16. The model is trained for 80 epochs. The initial learning rate is $2\times 10^{-4}$, and a cosine annealing strategy is used to gradually reduce the learning rate to $1\times 10^{-6}$. For the IVRG dataset, following the settings in [44], it is cropped into 16,374 training patches, and then additive Gaussian noise with random noise levels is used to simulate the target images, considering three noise levels in testing: $\sigma=25$, 50, and 75. The model is trained for 100 epochs with an initial learning rate of $5\times 10^{-4}$.
@@
Restormer [48] 31.72 0.876 27.67 0.788 24.50 0.722
Uformer [41] 31.97 0.878 27.83 0.793 24.58 0.726
MPRNet [47] 31.89 0.878 27.80 0.794 24.57 0.729
Xformer [49] 31.98 0.879 27.86 0.794 24.61 0.731
CUNet [11] 31.37 0.876 27.78 0.812 24.64 0.764
SANet [32] 31.54 0.885 28.13 0.825 24.81 0.785
MNNet [44] 32.59 0.904 28.66 0.851 25.25 0.808
DVN [18] 32.76 0.906 28.73 0.852 25.28 0.809
NIR-Restormer [43] 32.86 0.906 28.77 0.853 25.30 0.811
@@ 4.2 Experimental Results
Evaluations on the DVN dataset. We first evaluate our method on the DVD dataset, with quantitative results shown in Table 1. It can be observed that the proposed FCENet, and even FCENet-L, outperform all other single-image denoising methods and NIR-assisted denoising methods in terms of PSNR and SSIM with moderate computational and memory costs. The qualitative comparison in Fig. 6 clearly illustrates that our method has advantages over other state-of-the-art methods in noise removal, color preservation, and detail recovery. For single-image denoising methods like Restormer, although they have achieved advanced denoising effects in low-noise scenarios, they still lack effective means to address detail texture loss in high-noise situations. For guided denoising methods, MNNet is unable to effectively handle the inconsistencies between NIR and RGB images. DVN strengthens consistent high frequency textures but loses critical differential high frequency information in NIR and sacrifices the authenticity of low-frequency color. NIR-Restormer
Figure 8: Frequency domain dynamic filters on NIR and RGB side at different depths. The first row represents the frequency domain filters of RGB side, and the second row represents the filters of NIR side.
has good denoising effects but lacks effective utilization of inconsistent information, leading to over-smoothing and color distortion. Our method recovers detailed texture information while more faithfully preserving low-frequency color information in RGB images. This proves that the proposed method can efficiently utilize the complementary information of NIR and RGB images in the fusion through a series of frequency domain learning.
Evaluations on the IVRG dataset. We evaluate our method on the IVRG dataset. Following the previous method, we consider three noise levels in the test, including $\sigma=25,50$, and $75$. Quantitative results are shown in Table 2. It can be observed that compared to the state-of-the-art single-image denoising algorithms, our method significantly improves the performance of image denoising, demonstrating the effectiveness of NIR images. Compared to existing NIR-assisted denoising methods, our method also has significant advantages.
Evaluations on real-world experiment. To further evaluate the performance of our model in real noisy environments, we conduct qualitative experiments on NIR-RGB data pairs in real low-light conditions. As shown in Fig. 7, the introduction of NIR images brings more detailed textures but also a large number
@@
DFR 30.15 0.948
DFR + CFR 30.37 0.950
of inconsistencies. SANet fails to handle inconsistencies, resulting in a significant amount of artifacts, while MNNet, DVN and NIR-Restormer are unable to thoroughly denoise. In contrast, our method can eliminate inconsistencies while introducing more detailed information.
@@ 4.3 Ablation Study
We conduct ablation studies with FCENet on the DVD dataset under $\sigma=4$. We compare two mechanisms proposed in this paper with a two-stage U-Net baseline, with the results shown in Table 3. In the U-Net baseline, cross-field images are fused through addition. By deploying FDSM and FEFM, the model achieves gains of 0.51dB and 0.83dB, respectively. Finally, combining all contributions significantly improves the denoising level compared to the baseline.
Effect of FDSM. Using FDSM allows for better extraction of complementary information from NIR and RGB while handling inconsistent regions. In Fig. 8, we show the frequency domain dynamic filter kernels in FDSM. It can be observed that the filter kernels processing NIR feature maps regularly enhance the learning of mid-to-high frequency information, while those processing RGB feature maps mainly learn low frequency information from the input images. This is consistent with the conclusions drawn from our frequency domain correlation priors. Additionally, we conduct experiments using spatial dynamic convolution in this module. The results in Table 4 ultimately prove the effectiveness of extracting relevant features from the frequency domain in NIR and RGB images and the stability of dynamic convolution in extracting inconsistent features.
@@
Effect of FEFM. Traditional cross-attention fusion modules only strengthen common features but neglect differential features. FEFM strengthens the learning of common features and enhances the learning of some high frequency differential information in NIR feature maps. Fig. 9 shows the input NIR and RGB feature maps after frequency domain selection, the fused feature maps after cross-attention fusion, and the fused feature maps after FEFM. Compared to the NIR feature maps, the input RGB feature maps still lack high frequency features. It can be seen that the proposed FEFM strengthens common features while supplementing some differential texture information from NIR. We compare two mechanisms in FEFM, CFR and DFR, with traditional cross-attention fusion modules and direct feature summation. The results in Table 5 prove effectiveness of our method in strengthening common features in frequency domain and modeling differential features.
@@ 5 Conclusion
In this paper, a cross-field frequency correlation exploiting network for NIR-assisted image denoising is proposed. Specifically, we first conduct a series of analyses on the characteristics of NIR and RGB images in frequency domain and obtain frequency correlation prior. Based on the prior, the frequency dynamic selection mechanism dynamically extracts complementary information from cross-field images in the frequency domain, and the frequency exhaustive fusion mechanism is proposed to enhance common features and differential high frequency features during fusion process. Extensive experiments demonstrate that the proposed method is significantly better than other algorithms. It is worth mentioning that existing methods may fail on significantly misaligned cross-field images. In future work, we plan to explore advanced methods such as feature point matching or incorporating deformable convolutional kernels to handle significantly misaligned cross-field images.
Acknowledgments
This work is supported by the National Natural Science Foundation of China under Grants 62322204, 62131003, and the Key Laboratory of Target Cognition and Application Technology.